Early Payment discount decisions Prairie Manufacturing has four possible suppliers, all of which offer different credit termsplease use control scroll wheel to view the question if it is not visible.
Answer #1

a.
Cost of giving up cash discount
J : 2%/98% x (365/35) = 21.28%
K : 2%/98% x (365/70) = 10.64%
L : 1%/99% x (365/55) = 6.70%
M : 2%/98% x (365/90) = 8.28%

b.
The firm should give up L and M, since cost of giving up is less than bank rate

c.
There would be no difference to answer b, since cost of giving up was already lower than bank rate, stretching period with 30days will further reduce the cost of giving up